OpenStudy (anonymous):

Solving by addition x+y=4 x-y=6

OpenStudy (anonymous):

you got to be kidding

OpenStudy (anonymous):

what a jerk!

OpenStudy (saifoo.khan):

x+y=4 x-y=6 ===== 2x = 10 x = 10/2 x = 5 Now x = 5 in any equation u like to solve for y.

hero (hero):

Oh wow. Such strong emotions

OpenStudy (anonymous):

Not really that was just rude. it says y=-1 wheres that come from?

OpenStudy (anonymous):

once you have x=5 from adding the two equations together, you have to get the value for the y-coordnate by back substituting into any equation that has a y in it; i will show you using the first equation from the original system: x+y=4 5+y=4 (sub in the know value for x) y=-1 (subtract 5 from both sides of the equation) the answer may be written as x=5, y=-1, or as an ordered pair (5,-1) depending on the custom in your class.

OpenStudy (anonymous):

Thank you that was very helpful!

OpenStudy (anonymous):

ur welcome